# Defensive Model Expansion for Robust Bayesian Inference

### v1

@@ 2.3 Orthogonalization
When using models with predictions of the form $X_{i}^{\top}\beta+r(X_{i})$, such as Model NR or Model NC, it is helpful to require the nonparametric component $r(x)$ to be orthogonal to the linear predictors. That is, we require that $\sum_{i}r(X_{i})\,X_{i}=\mathbf{0}$. This condition serves two purposes. First, it can substantially improve the mixing of MCMC algorithms used to fit the model. Second, it prevents $r(\cdot)$ from becoming collinear with $X_{i}^{\top}\beta$, making it easier to shrink $r(\cdot)$ toward zero when it is unnecessary. Orthogonalization is not required for our theoretical results to hold, but it is often useful in practice. For GP models, orthogonalization is implemented by replacing the kernel $\kappa(x,x^{\prime})$ with the projected kernel $\kappa^{\star}(x,x^{\prime})=\kappa(x,x^{\prime})-k_{x}^{\top}\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{K}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top}k_{x^{\prime}},$ where $\boldsymbol{K}_{ij}=\kappa(X_{i},X_{j})$, $k_{x}=(\kappa(x,X_{1}),\ldots,\kappa(x,X_{N}))^{\top}$, and $\boldsymbol{X}\in\mathbb{R}^{N\times P}$ is the design matrix (Plumlee and Joseph,, 2018).
@@ 2.4 Summarizing Nonparametric Fits
One practical implementation is to form a linear projection summary by minimizing
$\text{SSE}=\sum_{i=1}^{N}\{\widehat{\mu}(X_{i})-X_{i}^{\top}\beta^{\star}\}^{2},$ (1)
where $\widehat{\mu}(x)$ is the posterior mean function (or a posterior draw). We then compute $R^{2}=1-\frac{\text{SSE}}{\sum_{i}\{\widehat{\mu}(X_{i})-\bar{\mu}\}^{2}},\text{ where }\bar{\mu}=N^{-1}\sum_{i}\mu(X_{i}).$ If $R^{2}$ is sufficiently close to 1, we conclude that the parametric model is adequate. The required closeness depends on context: if prediction is the main goal, we might demand $R^{2}$ very close to 1, while if the goal is only a broad scientific conclusion, a more modest $R^{2}$ could suffice. If the linear projection is inadequate, we expand to an additive model $\mu(x)\approx\sum_{j}\gamma_{j}(x_{j})$ and, if that is still insufficient, apply a CART fit to residuals to detect missing interaction terms. For logistic regression, it is natural instead to minimize the Kullback–Leibler divergence between the fitted probabilities $\mu(X_{i})$ and a logistic regression model with predictor $X_{i}^{\top}\beta^{\star}$.
Importantly, we do not recommend formal hypothesis testing of whether $r=r_{0}$. First, we may lack power to detect $r=r_{0}$ even if the model technically allows for it. Second, even when $r\neq r_{0}$ with high posterior probability, the parametric submodel may still be practically indistinguishable from the nonparametric model in terms of its predictions.
A criticism of the above strategy is that it is not specific enough, requiring subjective judgments about whether a summary is “adequate” or not. On the contrary, we feel that this is a strength, as it forces the analyst to engage with the question of what purpose the fitted model is supposed to serve and whether it is sufficient for that purpose. Judgments, such as what an appropriate tradeoff between interpretability and predictive accuracy are, should chiefly be determined in collaboration with subject matter experts.
@@ 3 Theoretical Results
We now establish some basic theoretical properties for the BART and GP models. To formalize the idea that a properly anchored Bayesian nonparametric model can perform “just as well” as a parametric submodel we will use arguments based on the rate of convergence of a posterior distribution; the rate will either be used as a justification on its own, or as part of the assumptions for stronger justifications. We will say that the posterior convergence rate is (upper bounded by) $\epsilon_{N}$ if we have $\Pi_{\alpha}\{H(\theta_{0},\theta)>M_{N}\,\epsilon_{N}\mid\boldsymbol{Z}_{N}\}\to 0$ in probability for all sequences $M_{N}\uparrow\infty$ (where $M_{N}$ is allowed to diverge arbitrarily slowly). Here, $H(\theta_{0},\theta)=[\int\{\sqrt{f_{\theta}(z)}-\sqrt{f_{\theta_{0}}(z)}\}^{2}\ dz]^{1/2}$ denotes the Hellinger distance between $f_{\theta}$ and $f_{\theta_{0}}$. For example, we will show that both GBART and the spike-and-GP approaches give $\epsilon_{N}=N^{-1/2}$ rates and that posterior inferences for $\beta_{0}$ are efficient, provided that the parametric submodel is correct.
@@
where $L_{N}(\theta)=\prod_{i=1}^{N}f_{\theta}(Z_{i})$ is the likelihood. Works that have studied the properties of fractional posteriors include O’Hagan, (1995); Walker and Hjort, (2001); Bhattacharya et al., (2019); Martin and Walker, (2019). Remarkably, under essentially no further assumptions, the prior thickness condition is sufficient to guarantee posterior concentration at the rate $\epsilon_{N}$. Section 2.2 of Castillo, (2024) is particularly instructive for understanding why these conditions are sufficient and how the arguments can be extended to non-iid settings.
@@
The role of $\alpha$ can be understood in two ways. First, it is as if we “hold out” $100(1-\alpha)\%$ of the data, which makes the posterior less prone to overfitting (O’Hagan,, 1995). Second, as noted by Walker and Hjort, (2001), the fractional posterior is equivalent to using the full likelihood $L_{N}(\theta)$ with a data-adaptive prior $\Pi^{\star}(d\theta)\propto\Pi(d\theta)/L_{N}(\theta)^{1-\alpha}$. This construction downweights parameter values that fit the observed data too well, again limiting overfitting. This robustness to overfitting is what allows us to avoid conditions on the size of the effective support required of the full posterior.
Identifiability
Prior thickness ensures convergence in Hellinger distance, but does not automatically imply efficient estimation of $\eta_{0}$, since $r(\cdot)$ may be confounded with $\eta$. For example, in Models NR and NC, $\mu(\cdot)\to\mu_{0}(\cdot)$ does not necessarily imply $\beta\to\beta_{0}$. One way to ensure recovery of $\eta_{0}$ is through a metric identifiability condition such as $\|\eta-\eta_{0}\|\lesssim H(\theta,\theta_{0}),$ which forces $H(\theta,\theta_{0})\to 0$ only if $\eta\to\eta_{0}$ at the same rate. In Model NR this can be obtained by imposing $\mathbb{E}_{\theta}\{X_{i}r(X_{i})\}=\mathbf{0}$ and assuming $\mathbb{E}(X_{i}\,X_{i}^{\top})=\Sigma_{X}$ is full-rank. Alternatively, consistent selection of the event $r=r_{0}]$ also guarantees efficient inference for $\eta_{0}$.
@@
It is possible to extend our results to allow for non-iid data by instead defining $B_{\epsilon}$ in terms of $\mathbb{E}_{\theta_{0}}\left(\log\frac{f(Z_{1},\ldots,Z_{N}\mid\theta_{0})}{f(Z_{1},\ldots,Z_{N}\mid\theta)}\right)^{k}\leq N\,\epsilon_{N}^{2}$ for $k\in\{1,2\}$, and similarly redefining the Hellinger distance. It is also possible to convert the convergence rate arguments into non-asymptotic arguments. Finite sample results are given, for example, by Bhattacharya et al., (2019).
@@
The usual posterior ($\alpha=1$) requires additional technical work that we feel distracts from the message of this paper. In this case, prior thickness alone is not enough; one must also control the complexity of the support of the prior, often by placing high mass on a low-entropy sieve (Ghosal et al.,, 2000). Prior support conditions serve to stop overfitting to the data. While this strategy works, it places much stronger restrictions on the choice of $\Pi_{r}(dr)$ and can introduce an extra $(\log N)^{t}$ factor in the rates that can be cumbersome to remove.
@@ 3.1 General Results
We start with general theoretical results that will later be applied to BART and GP regression models. Throughout this section (and Sections 3.2 and 3.3) we fix $\alpha\in(0,1)$. Our first result is a general fractional posterior concentration theorem, implied for example by Theorem 2.1 of Castillo, (2024).
@@
Suppose that PPA holds and that for all sequences $M_{N}\uparrow\infty$ there exist positive constants $C_{1},C_{2},N^{\star}$ such that $\Pi_{r}(\|r-r_{0}\|_{\mathcal{H}}\leq M_{N}N^{-1/2})\geq C_{1}\,e^{-C_{2}\,M_{N}^{2}}$ for $N\geq N^{\star}$. Then for all sequences $M_{N}\uparrow\infty$ we have $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}\{H(\theta_{0},\theta)>M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ as $N\to\infty$.
@@
Suppose that PPA holds and that $\Pi_{r}(r=r_{0})>0$. Then for all sequences $M_{N}\uparrow\infty$ we have $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}\{H(\theta_{0},\theta)\geq M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ as $N\to\infty$.
Anchoring to the parametric model also enables consistent model selection. If the parametric submodel holds and the nonparametric part cannot achieve the parametric rate on its own, then the posterior concentrates all its mass on the parametric model.
@@
and suppose that $\mathbb{E}_{\theta_{0}}\Pi^{\star}_{\alpha}\{H(\theta_{0},\theta)<M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ as $N\to\infty$. Then $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}(r=r_{0}\mid\boldsymbol{Z}_{N})\to 1$ as $N\to\infty$.
@@
The assumption that $\Pi^{\star}_{\alpha}(\cdot\mid\boldsymbol{Z}_{N})$ fails to achieve the parametric rate is stronger than strictly necessary, as fractional Bayes factors are typically consistent for nested parametric model selection (O’Hagan,, 1995). This assumption is not very restrictive, however, because truly nonparametric models rarely achieve parametric rates unless they place positive mass on a correctly specified submodel.
When Theorem 2 is combined with a parametric submodel $\mathcal{F}_{0}$ that is differentiable in quadratic mean (see Section 5.5 of van der Vaart,, 1998), we also obtain a Bernstein–von Mises result. The following theorem formalizes this under standard regularity conditions.
@@
The value of Theorem 3 relative to the usual Bernstein-von Mises theorem is that it represents an oracle property of the fractional posterior: if the parametric model is correctly specified, our nonparametric prior will nevertheless obtain efficient inference for $\eta_{0}$. On the other hand, if the parametric model is misspecified we will also be able to obtain efficient nonparametric estimation rates for $\theta_{0}$; the exact rates we get under misspecification will depend on the choice of $\Pi^{\star}(dr)$ and on $r_{0}(x)$.
@@ 3.2 Results for Bayesian Additive Regression Trees
Then the conditions of Theorem 1 hold. Additionally, let $\mathcal{N}$ be the event that all decision trees in the ensemble contain no splitting rules. Then $\Pi_{\alpha}(\mathcal{N}\mid\boldsymbol{Z}_{N})\to 1$ in $F_{\theta_{0}}$-probability.
This result is close in spirit to Theorem 2. While we cannot directly apply Theorem 3, Theorem 4 is strong enough to imply a Bernstein–von Mises result.
@@
Suppose that the conditions of Theorem 4 hold and that $F_{X}$ is such that $X_{i1}\equiv 1$ and that $\mathbb{E}(X_{i}\,X_{i}^{\top})=\Sigma_{X}$ is non-singular. Define the vector $\eta=(\beta_{1}+r(0),\beta_{2},\beta_{3},\ldots,\beta_{D},\sigma)^{\top}$ and let $\widehat{\eta}$ by the maximum likelihood estimator of $(\beta,\sigma)$ in Model NR when $r_{0}\equiv 0$. Then the conclusion of Theorem 3 holds.
Remark 2.
Versions of the above results can also be derived for Model NC and Model DR; we only give results for Model NR for brevity.
3.3 Results for Gaussian Processes
@@
Applying Theorem 2 directly requires us to have a lower bound on the rate of convergence for Gaussian process models. In other words, we need to verify that a pure GP prior without a spike-and-slab component does not achieve the parametric $N^{-1/2}$ rate. To keep technical details to a minimum, we find it prudent to adopt $\Pi_{\alpha}\{H(\theta_{0},\theta)>M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ in $F_{\theta_{0}}$-probability as a starting assumption rather than to state conditions for it to hold. This will almost always be the case when $\kappa(x,x^{\prime})$ is a strictly positive definite kernel such as the squared exponential and Matérn kernels.
@@
We might expect flexible nonparametric methods to attain a rate of the form $\epsilon_{N}=N^{-p/(2p+D)}$ for some $p>0$. For some common kernels (such as the squared exponential kernel), however, we only miss $\epsilon_{N}=N^{-1/2}$ by a logarithmic factor (see Theorem 10 of van der Vaart and van Zanten,, 2011). This occurs because the reproducing kernel Hilbert space (RKHS) of the squared exponential kernel consists of extremely smooth functions, making it particularly well suited to approximating the function $r_{0}(x)=0$. The general problem of lower bounding posterior contraction rates for Gaussian process regression models is studied by Castillo, (2008).
If the goal is inference on $\beta_{0}$, it can also suffice to not apply shrinkage and instead project the GP posterior onto the family $\mathcal{F}_{0}$. Recall that using projections for inference is what we actually recommend. For Model NR this is particularly simple due to conjugacy.
@@
For a Gaussian process prior with Model NR, the posterior distribution of $\beta^{\star}$ is normal, and so Theorem 6 really only concerns the behavior of the posterior mean and variance. We note that Theorem 6 does not directly rule out that a Bernstein-von Mises result holds without the linear term $\sigma^{2}_{\beta}x^{\top}x^{\prime}$ in the kernel. In Section 4.3 we confirm empirically that Theorem 6 accurately predicts performance, while kernels without the linear term are empirically very highly biased in finite samples.
Theorem 6 is proven using the conjugacy of the GP prior. To show that the result is general, we also prove an analogous result for logistic regression (Model NC). The proof follows Theorem 2.2 of L’Huillier et al., (2023), which is itself based on Theorem 2.1 of Castillo and Rousseau, (2015). The main ingredients of the proof are that (i) $\|r-r_{0}\|=o(N^{-1/4})$ with high posterior probability, which allows us to control the remainder of a Taylor expansion and (ii) $r_{0}(x)=x^{\top}\beta_{0}$ lies in the RKHS of the prior because of the linear term in the kernel.
Theorem 7 (Logistic GP Bernstein–von Mises).
Consider Model NC with $r_{0}(x)\equiv 0$ and $X_{1},X_{2},\ldots$ a deterministic sequence in $0,1]^{P}$. Suppose $\widehat{I}_{N}=N^{-1}\sum_{i}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}X_{i}X_{i}^{\top}$ is invertible and $\widehat{I}_{N}\to I$ for some positive-definite $P\times P$ matrix $I$. Let $\Pi(\cdot\mid\boldsymbol{Z}_{N})$ be the posterior distribution of $r(\cdot)$ associate to the model with $Y_{i}\mid X_{i},r]\sim\operatorname{Bernoulli}\{(1+e^{-r(X_{i})})^{-1}\}$ and $r\sim\operatorname{GP}(0,\kappa)$ where $\kappa(x,x^{\prime})=\sigma_{\beta}^{2}x^{\top}x^{\prime}+\kappa^{\star}(x,x^{\prime})$ for $\sigma_{\beta}^{2}>0$ and $\kappa^{\star}$ satisfying Assumption A in the Supplementary Material. Define $\beta^{\star}=\widehat{I}_{N}^{-1}N^{-1}\sum_{i}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}X_{i}r(X_{i})$. Then
$d\left\{\sqrt{N}(\beta^{\star}-\widehat{\beta}),\operatorname{Normal}(0,I^{-1})\right\}\to 0\quad\text{in $F_{\theta_{0}}$-probability}$

### v2

@@ 2.3 Orthogonalization
When using models with predictions of the form $X_{i}^{\top}\beta+r(X_{i})$, such as Model NR or Model NC, it is helpful to require the nonparametric component $r(x)$ to be orthogonal to the linear predictors. That is, we require that $\sum_{i}r(X_{i})\,X_{i}=\mathbf{0}$. Here this is meant primarily as an empirical, finite-sample orthogonalization device based on the observed design matrix rather than as a population-level identifiability assumption. This condition serves two purposes. First, it can substantially improve the mixing of MCMC algorithms used to fit the model. Second, it prevents $r(\cdot)$ from becoming colinear with $X_{i}^{\top}\beta$, making it easier to shrink $r(\cdot)$ toward zero when it is unnecessary. Orthogonalization is not required for our theoretical results to hold, but it is often useful in practice. For GP models, orthogonalization is implemented by replacing the kernel $\kappa(x,x^{\prime})$ with the projected kernel $\kappa^{\star}(x,x^{\prime})=\kappa(x,x^{\prime})-k_{x}^{\top}\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{K}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top}k_{x^{\prime}},$ where $\boldsymbol{K}_{ij}=\kappa(X_{i},X_{j})$, $k_{x}=(\kappa(x,X_{1}),\ldots,\kappa(x,X_{N}))^{\top}$, and $\boldsymbol{X}\in\mathbb{R}^{N\times P}$ is the design matrix (Plumlee and Joseph,, 2018).
@@ 2.4 Summarizing Nonparametric Fits
One practical implementation is to form a linear projection summary by minimizing $\text{SSE}=\sum_{i=1}^{N}\{\mu(X_{i})-X_{i}^{\top}\beta^{\star}\}^{2}.$ We then compute $R^{2}=1-\frac{\text{SSE}}{\sum_{i}\{\mu(X_{i})-\bar{\mu}\}^{2}},\text{ where }\bar{\mu}=N^{-1}\sum_{i}\mu(X_{i}).$ We can interpret $R^{2}$ as the proportion of the overall signal that is explained by the parametric model. If $R^{2}$ is sufficiently close to 1, we conclude that the parametric model is adequate. The required closeness depends on context: if prediction is the main goal, we might demand $R^{2}$ very close to 1, while if the goal is only a broad scientific conclusion, a more modest $R^{2}$ could suffice. As a reporting convention, we recommend presenting the posterior mean and posterior distribution of $R^{2}$; for many descriptive purposes, values above roughly $0.99$ indicate an essentially interchangeable summary, values around $0.95$ indicate a close but imperfect summary, and materially smaller values suggest that richer structure may matter. If the linear projection is inadequate, we expand to an additive model $\mu(x)\approx\sum_{j}\gamma_{j}(x_{j})$ and, if that is still insufficient, apply a CART fit to residuals to detect missing interaction terms. For logistic regression, it is natural instead to minimize the Kullback–Leibler divergence between the fitted probabilities $\mu(X_{i})$ and a logistic regression model with predictor $X_{i}^{\top}\beta^{\star}$.
Importantly, we do not recommend formal hypothesis testing of whether $r=r_{0}$. First, we may lack power to detect $r=r_{0}$ even if the model technically allows for it. Second, even when $r\neq r_{0}$ with high posterior probability, the parametric submodel may still be practically indistinguishable from the nonparametric model in terms of its predictions; this can occur in the GBART model, for example, when a shrinkage prior is placed on $\sigma_{\mu}$.
A criticism of the above strategy is that it is not specific enough, requiring subjective judgments about whether a summary is “adequate” or not. On the contrary, we feel that this is a strength, as it forces the analyst to engage with the question of what purpose the fitted model is supposed to serve and whether it is sufficient for that purpose. Judgments, such as what an appropriate tradeoff between interpretability and predictive accuracy are, should chiefly be determined in collaboration with stakeholders.
@@ 3 Theoretical Results
We now establish some basic theoretical properties for the BART and GP models. The results in this section split into two parts: general fractional-posterior concentration and model-selection results under correct parametric specification, and ordinary-posterior semiparametric Bernstein-von Mises results for projection parameters. To formalize the idea that a properly anchored Bayesian nonparametric model can perform “just as well” as a parametric submodel, we will use arguments based on the rate of convergence of a posterior distribution; the rate can be used as a justification on its own, or as part of the assumptions for stronger justifications such as model selection consistency or a Bernstein-von Mises theorem. We will say that the posterior convergence rate is (upper bounded by) $\epsilon_{N}$ if we have $\Pi_{\alpha}\{H(\theta_{0},\theta)>M\,\epsilon_{N}\mid\boldsymbol{Z}_{N}\}\to 0$ in probability for some $M>0$. Here, $H(\theta_{0},\theta)=[\int\{\sqrt{f_{\theta}(z)}-\sqrt{f_{\theta_{0}}(z)}\}^{2}\ dz]^{1/2}$ denotes the Hellinger distance between $f_{\theta}$ and $f_{\theta_{0}}$. For example, we will show that appropriate GBART and spike-and-GP priors give $\epsilon_{N}=N^{-1/2}$ rates (possibly up-to a logarithmic term) and that posterior inferences for $\beta_{0}$ are efficient, provided that the parametric submodel is correct.
@@
where $L_{N}(\theta)=\prod_{i=1}^{N}f_{\theta}(Z_{i})$ is the likelihood; extending to $\alpha=1$ (i.e., the genuine posterior) is possible for the models we consider here under additional conditions on the prior. Works that have studied the properties of fractional posteriors include O’Hagan, (1995); Walker and Hjort, (2001); Bhattacharya et al., (2019); Martin and Walker, (2019). Under essentially no further assumptions, the prior thickness condition is sufficient to guarantee posterior concentration at the rate $\epsilon_{N}$. Section 2.2 of Castillo, (2024) is particularly instructive for understanding why these conditions are sufficient and how the arguments can be extended to non-iid settings.
@@
The role of $\alpha$ can be understood in two ways. First, it is as if we “hold out” $100(1-\alpha)\%$ of the data, which makes the posterior less prone to overfitting (O’Hagan,, 1995). Second, as noted by Walker and Hjort, (2001), the fractional posterior is equivalent to using the full likelihood $L_{N}(\theta)$ with a data-adaptive prior $\Pi^{\star}(d\theta)\propto\Pi(d\theta)/L_{N}(\theta)^{1-\alpha}$. This construction downweights parameter values that fit the observed data too well, again limiting overfitting. This robustness to overfitting is what allows us to avoid conditions on the size of the effective support used, for example, by Ghosal et al., (2000) to study the usual posterior.
@@
It is possible to extend our results to allow for non-iid data by instead defining $B_{\epsilon}$ in terms of $\mathbb{E}_{\theta_{0}}\left(\log\frac{f(Z_{1},\ldots,Z_{N}\mid\theta_{0})}{f(Z_{1},\ldots,Z_{N}\mid\theta)}\right)^{k}\leq N\,\epsilon_{N}^{2}$ for $k\in\{1,2\}$, and similarly redefining the Hellinger distance. Another advantage of fractional posteriors is that it is also possible to convert convergence rate arguments into non-asymptotic arguments and allow for model misspecification, as done by Bhattacharya et al., (2019).
@@
The usual posterior ($\alpha=1$) requires additional technical work that we feel distracts from the message of this paper. In this case, prior thickness alone is not enough; one must also control the complexity of the support of the prior, often by placing high mass on a low-entropy sieve (Ghosal et al.,, 2000). While this strategy works, it places much stronger restrictions on the choice of $\Pi_{r}(dr)$. Nevertheless, we note that the specific semiparametric Bernstein-von Mises results given in Theorem 5, Theorem 6, and Theorem 7 happen to be true for genuine posteriors; the fractional posterior only makes establishing the required rate of convergence easier, but convergence rate results for the non-fractional posterior are widely available for BART and GP models.
@@ 3.1 General Results
We start with general theoretical results that will later be applied to BART and GP regression models. Throughout this section (and Sections 3.2 and 3.3) we fix $\alpha\in(0,1)$. Our first result is a general fractional posterior concentration theorem, implied for example by Theorem 2.1 of Castillo, (2024). It establishes a near-parametric rate of convergence under essentially only a prior thickness condition.
@@
Suppose that PPA holds and that there exist constants $C_{1},C_{2},N^{\star}$ such that $\Pi_{r}(\|r-r_{0}\|_{\mathcal{H}}\leq M_{N}N^{-1/2})\geq C_{1}e^{-C_{2}M_{N}^{2}}$ where $M_{N}/\sqrt{\log N}\to\infty$. Then we have $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}\{H(\theta_{0},\theta)>K\,M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ as $N\to\infty$ for some sufficiently large $K$.
@@
Suppose that PPA holds, that $\Pi_{r}(r=r_{0})>0$, and that $M_{N}/\sqrt{\log N}\to\infty$. Then we have $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}\{H(\theta_{0},\theta)\geq M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ as $N\to\infty$.
Anchoring to the parametric model also enables consistent model selection. If the parametric submodel holds and the nonparametric part cannot achieve a near-parametric rate on its own, then the posterior concentrates all its mass on the parametric model.
@@
and suppose that $\mathbb{E}_{\theta_{0}}\Pi^{\star}_{\alpha}\{H(\theta_{0},\theta)<M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ where $M_{N}/\sqrt{\log N}\to\infty$ as $N\to\infty$. Then $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}(r=r_{0}\mid\boldsymbol{Z}_{N})\to 1$ as $N\to\infty$.
@@
The assumption that $\Pi^{\star}_{\alpha}(\cdot\mid\boldsymbol{Z}_{N})$ fails to achieve a near-parametric rate is stronger than strictly necessary, as fractional Bayes factors are typically consistent for nested parametric model selection (O’Hagan,, 1995) (where the rate is $N^{-1/2}$ for both models). We discuss settings where it is obtained for GP priors in Section 3.3, where we note that some GP priors can obtain near-parametric rates; these models fit more naturally with the posterior projection approach described in Section 3.3 in terms of analysis, as a near-parametric rate is useful for proving a Bernstein-von Mises result.
When Theorem 2 is combined with a parametric submodel $\mathcal{F}_{0}$ that is differentiable in quadratic mean (see Section 5.5 of van der Vaart,, 1998), we also obtain a Bernstein-von Mises result. The following theorem formalizes this under standard regularity conditions.
@@
The value of Theorem 3 relative to the usual Bernstein-von Mises theorem is that it represents an oracle property of the fractional posterior: if the parametric model is correctly specified, our nonparametric prior will nevertheless obtain efficient inference for $\eta_{0}$. Efficiency holds up-to inflation of the variance by a factor of $\alpha^{-1}$; this can be corrected using the shift-and-rescale approach of L’Huillier et al., (2023). On the other hand, if the parametric model is misspecified we will also be able to obtain efficient nonparametric estimation rates for $\theta_{0}$; the exact rates we get under misspecification will depend on the choice of $\Pi^{\star}(dr)$ and on $r_{0}(x)$.
@@ 3.2 Results for Bayesian Additive Regression Trees
Then the conditions of Theorem 1 hold. Additionally, let $\mathcal{N}$ be the event that all decision trees in the ensemble contain no splitting rules. Then $\mathbb{E}_{0}\Pi_{\alpha}(\mathcal{N}\mid\boldsymbol{Z}_{N})\to 1$.
This result is close in spirit to Theorem 2. While we cannot directly apply Theorem 3, Theorem 4 is strong enough to imply a Bernstein-von Mises result.
@@
Suppose that the conditions of Theorem 4 hold and that $F_{X}$ is such that $X_{i1}\equiv 1$ and that $\mathbb{E}(X_{i}\,X_{i}^{\top})=\Sigma_{X}$ is non-singular. Define the vector $\eta=(\beta_{1}+r(0),\beta_{2},\beta_{3},\ldots,\beta_{D},\sigma)^{\top}$ and let $\widehat{\eta}$ be the maximum likelihood estimator of $(\beta,\sigma)$ in Model NR when $r_{0}\equiv 0$. Then the conclusion of Theorem 3 holds.
3.3 Results for Gaussian Processes and Posterior Projections
@@
Applying Theorem 2 directly requires us to have a lower bound on the rate of convergence for Gaussian process models. In other words, we need to verify that a GP prior without a spike-and-slab component does not achieve the near-parametric $(\log N/N)^{1/2}$ rate. We might expect flexible nonparametric methods to attain a rate of the form $\epsilon_{N}=N^{-p/(2p+D)}$ up-to logarithmic terms for some $p>0$; this is the case for Matérn priors (see see Theorem 5 of van der Vaart and van Zanten,, 2011 and the discussion thereafter). For some common kernels (such as the squared exponential kernel), however, a near-parametric rate is attained (see Theorem 10 of van der Vaart and van Zanten,, 2011). This occurs because the squared exponential kernel is supported on extremely smooth functions, making it particularly well suited to approximating the function $r_{0}(x)=0$. The general problem of lower bounding posterior contraction rates for Gaussian process regression models is studied by Castillo, (2008).
As an alternative to establishing consistent model selection, if the goal is inference on $\beta_{0}$ it can also suffice to instead project the posterior of $(r,\beta)$ onto the family $\mathcal{F}_{0}$. Recall that using projections for inference is what we actually recommend. The following two theorems give results for posterior projections of models based on Gaussian process priors; we state both results for the non-fractional posterior.
@@
Xie and Xu, (2021) study Model NR in detail, and also allow for $x^{\top}\beta$ to be replaced by a nonlinear term $g(x;\beta)$ under appropriate conditions. Additionally, they allow $r_{0}(x)\neq 0$ and impose only that $r_{0}(x)$ is sufficiently smooth, with $\Sigma_{X}^{-1}$ replaced with the semiparametric efficient variance in Theorem 6.
Remark 4.
For a Gaussian process prior with Model NR, the posterior distribution of $\beta^{\star}$ is normal, and so Theorem 6 really only concerns the behavior of the posterior mean and variance. We note that Theorem 6 does not directly rule out that a Bernstein-von Mises result holds without the linear term $\sigma^{2}_{\beta}x^{\top}x^{\prime}$ in the kernel. In Section 4.3 we confirm empirically that Theorem 6 accurately predicts performance, while kernels without the linear term are empirically highly biased in finite samples.
To show that this phenomenon is general, we prove an analogous result for logistic regression (Model NC). The proof follows Theorem 2.2 of L’Huillier et al., (2023), which is itself based on Theorem 2.1 of Castillo and Rousseau, (2015). The main steps of the proof are (i) we assume that $\|r-r_{0}\|_{N}=\sqrt{N^{-1}\sum_{i}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}\{r_{0}(X_{i})-r(X_{i})\}^{2}}$ has a posterior contraction rate faster than $N^{-1/4}$, which allows us to control the remainder of a Taylor expansion and (ii) a change-of-variables argument under the prior $\beta\sim\operatorname{Normal}(\mu_{\beta},\Sigma_{\beta})$, similar to the application of the Cameron-Martin theorem by Castillo and Rousseau, (2015). On the technical side, guaranteeing that item (i) holds is difficult due to the fact that $\|r-r_{0}\|_{N}$ is not equivalent to the empirical Hellinger distance, and so is not available from existing results for Gaussian process logistic regression; as discussed in the Supplementary Material, this difficulty can be bypassed by working with an appropriately-truncated prior instead. As with the results of Xie and Xu, (2021), we expect that this result can be modified with additional assumptions to also accommodate $r_{0}(x)$ deviating from $x^{\top}\beta_{0}$ provided that $\|r-r_{0}\|_{N}$ has a posterior contraction rate faster than $N^{-1/4}$; after such a modification, the asymptotic variance would be the semiparametric efficient variance.
Theorem 7 (Bernstein-von Mises for Logistic Regression).
Consider Model NC with $r_{0}(x)\equiv 0$ and consider a fixed sequence $X_{1},X_{2},\ldots$ in $0,1]^{P}$ and let $\widehat{I}_{N}=\frac{1}{N}\sum_{i=1}^{N}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}X_{i}\,X_{i}^{\top}$. Suppose $\widehat{I}_{N}$ is invertible and that $\widehat{I}_{N}\to I$ for some positive definite matrix $I$. Let $\Pi(\cdot\mid\boldsymbol{Z}_{N})$ be the posterior distribution of $(r,\beta)$ associated to the model with $Y_{i}\mid X_{i},r,\beta]\sim\operatorname{Bernoulli}\{(1+e^{-(X_{i}^{\top}\beta+r(X_{i}))})^{-1}\}$ such that Assumption A in the Supplementary Material is satisfied. Define $\beta^{\star}=\widehat{I}_{N}^{-1}N^{-1}\sum_{i}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}X_{i}\{r(X_{i})+X_{i}^{\top}\beta\}$ where $\widehat{I}_{N}=N^{-1}\sum_{i}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}X_{i}\,X_{i}^{\top}$. Then
$d\left\{\Pi\left(\sqrt{N}(\beta^{\star}-\widehat{\beta})\in\cdot\mid\boldsymbol{Z}_{N}\right),\operatorname{Normal}(0,I^{-1})\right\}\to 0\quad\text{in $F_{\theta_{0}}$-probability}$